# VSF: Simple, Efficient, and Effective Negative Guidance in Few-Step Image Generation Models By Value Sign Flip

### v2

@@ 4.9 Abliation Study
Figure 9: Trade Off Plot For Abliation Study
Figure 10: Example of Whole Embedding Flip (WEF), where the negative prompt embedding got flipped and concated with the positive prompt embedding before sending into a normal DiT
@@
Rather than altering the attention values, we explored a simpler and more intuitive approach: flipping the text embedding prior to input into the DiT (Whole Embedding Flip, WEF). This is similar to applying the CFG on text embeddings in NASA, but keeps the positive and negative tokens separated. Specifically, the negative text embedding is scaled by $\alpha$, concatenated with the positive prompt embedding in the sequence length dimension, and used as the prompt embedding for the DiT. We did not remove the padding for the negative prompt, as we found out that removing it causes the negative prompts to have no effects at all. We also tested our approach with no bias, no mask (but still duplication), and no duplication, no mask. The trade-off plot is shown in Figure 9. We specifically show the results from WEF in Figure 10 across different scales. This simpler and more intuitive approach appears to have no effect. We hypothesize that this is because it is similar to flipping both the key and the value, causing regions most similar to the flipped key (i.e., least similar to the original negative prompt) to be pushed away, rather than pushing away regions most similar to the original negative prompt (i.e., unflipped key). From the figure, we can see that the configurations without masking have a sharp positive score drop as the negative score increases. The WEF has a very limited range of negative scores, confirming the qualitative results. Our methods and the one without attention bias has similar results, showing that the bias is optional and can be compensated with scaling changes.
Positive Score Negative Score Quality Score
VSF 0.870 0.545 0.952
NAG (Chen et al., 2025a) 0.993 0.220 ($p<10^{-6}$) 0.968
NAG++ 0.975 0.320 ($p<10^{-6}$) 0.901
NASA(Nguyen et al., 2024) 0.970 0.380 ($p<10^{-6}$) 0.867
None 0.990 0.195 ($p<10^{-6}$) 0.968
CFG (Ho & Salimans, 2022) (non-few-step) 1.000 0.300 ($p<10^{-6}$) 0.956
@@
Table 2: Positive scores (how well the model follows the positive prompts) and negative scores (how well the model avoids the negative prompts) of our model (VSF), NAG (Chen et al., 2025a), and NAG with hyperparameter re-tuned (NAG++).
4.10 Adapting to Other DiT Models
@@
In this paper, we primarily use SD-3.5 (sd, 3) due to simplicity and elegant architecture. However, our method can theoretically be adapted to any transformer-based diffusion or flow-matching model. To demonstrate this adaptability, we implemented our method on Wan 2.1 with CausVid LoRA (Yin et al., 2025; Seppanen, ).
@@
For Wan 2.1, which uses cross-attention between image and text, masking is unnecessary and not used. Because our approach does not perform extrapolation and solely provides negative guidance, it cannot enhance overall quality significantly or replace CFG sampling in non-disstilled models, making it incompatible with the original Wan 2.1 model. Instead, we utilize CausVid (Yin et al., 2025), which enables Wan to function effectively without classifier-free guidance in few-step settings. Specifically, we used a LoRA (Seppanen, ) distilled from the original CausVid that can be directly applied on top of Wan 2.1. For qualitative results from Wan, please see the appendix.
We also tested our method on Flux Schnell (Black Forest Lab, 2025). However, due to its architecture combining one-stream and two-stream attention mechanisms, our approach did not significantly impact its tendency to ignore negative prompts. Future work should investigate these differences and explore ways to improve effectiveness.
4.11 Computational Cost
@@
Since our method does not require two passes through the entire model (as in CFG) or the attention module (as in NAG or NASA), and only slightly increases the sequence length, its theoretical computatio nal cost is significantly lower, close to that of a single pass. However, due to implementation limitations (specifically, FlashAttention-2’s lack of support for arbitrary attention masking), the actual runtime of our method is higher than the original single-pass MM-DiT models, and similar to NAG or NASA but much lower than CFG.
@@
To accurately measure the computational cost, we evaluate the runtime of 25 identical prompts under four settings: no guidance, NAG, NASA, and our proposed guidance, VSF, and then report the average runtime and peak memory usage for each setting. To avoid GPU thermal throttling affecting the results, we pause for at least 5 minutes between each set of tests. The tests are done on NVIDIA A100 40GB on Google Colab, as this is the most accessible option for high-end GPUs for users. Stable Diffusion Turbo is generated in 8 steps for 1024x1024 resolution, Wan is generated in 8 steps with 480x832 resolution, and 81 frames. The results are shown in Table 3.

### v5

@@ 6.3 Abliation Study
Rather than altering the attention values, we explored a simpler and more intuitive approach: flipping the text embedding prior to input into the DiT (Whole Embedding Flip, WEF). This is similar to applying the CFG on text embeddings studied in Nguyen et al. (2024), but keeps the positive and negative tokens separated. Specifically, the negative text embedding is scaled by $\alpha$, concatenated with the positive prompt embedding in the sequence length dimension, and used as the prompt embedding for the DiT. We did not remove the padding for the negative prompt, as we found out that removing it causes the negative prompts to have no effect at all.
We also tested our approach with no bias, no mask (but still duplication), and no duplication no mask. The trade-off plot is shown in Figure 6 Right. The simpler and more intuitive WEF approach appears to have almost no effect at all. We hypothesize that this is because it is similar to flipping both the key and the value, causing regions most similar to the flipped key (i.e., least similar to the original negative prompt) to be pushed away, rather than pushing away regions most similar to the original negative prompt (i.e., unflipped key). From the figure, we can see that the configurations without masking have a sharp positive score drop as the negative score increases. The WEF has a very limited range of negative scores. Our methods and the one without attention bias have similar results; this could be due to the MLLM not being sensitive enough to see the minor changes in quality.
Ablation study on hyperparameters is shown in the Appendix.
@@
Table 2: Positive scores (how well the model follows the positive prompts) and negative scores (how well the model avoids the negative prompts) of our model (VSF), NAG (Chen et al., 2025a), and NAG with hyperparameter re-tuned (NAG Strong).
Positive Score ($\uparrow$) Negative Score ($\uparrow$) Quality Score ($\uparrow$)
VSF Strong 0.870 0.545 0.952
VSF Quality 0.980 0.420 0.986
NAG (Chen et al., 2025a) 0.993 0.220 0.968
NAG Strong 0.975 0.320 0.901
NASA(Nguyen et al., 2024) 0.970 0.380 0.867
None 0.990 0.195 0.968
CFG (Ho & Salimans, 2022) (28 steps) 1.000 0.300 0.956
Table 3: Human Labelled Metric For 10 Selected Prompts with 2 Seeds
· Positive Score ($\uparrow$) Negative Score ($\uparrow$) Quality Score ($\uparrow$) NAG Strong 0.950 0.250 0.675 NAG 1.000 0.100 0.895 NASA 0.950 0.150 0.685 VSF Quality 0.900 0.550 0.823
7 Conclusion
In this paper, we introduced VSF, a novel approach for enhancing negative prompt adherence in image and video generation models. Our method involves flipping the sign of attention values and duplicating negative prompts and attention masking, effectively suppressing unwanted content. Experimental results indicate that VSF significantly outperforms previous methods, NAG Chen et al. (2025a), NASA (Nguyen et al., 2024), and even CFG in terms of negative prompt adherence, with much lower trade-offs in overall quality and positive prompt following. We also showed that VSF can be applied to create more creative (by style avoidance, abstract images, and anti-aesthetics styles) images. VSF also only has one main hyperparameter and one minor hyperparameter, making it easier to tune them in downstream tasks. Future work directions are discussed in the Appendix.
8 Reproducibility Statement
All code, dataset, and fine-tuned models (NegAwareQwen) will be released after publication.
Acknowledgements
Appendix
Appendix
Appendix A Negation in Vision Language Models
Much previous work has shown that existing vision language models (VLM) struggle to understand negation (Yuksekgonul et al., 2023; Singh et al., 2024; Alhamoud et al., 2025; Park et al., 2025). In classification tasks, the model cannot correctly understand text with negation in it, e.g. “a dog running” vs “a dog not running” might have very close embeddings, even though they are opposite. In our test using a CLIP-ViT-Base-32, the cosine similarity between “a dog running” and “a dog not running” is 0.9243, where the similarity between “a dog” and “a dog running” is only 0.8710. In Figure 7, we show 4 prompts “a photo of a bike”, “a photo of a bike without wheels”, “a photo of a bike with wheels”, and “a photo of a car with wheels”. We can see that the bike with wheels and the bike without wheels have extremely close embeddings.
Figure 7: PCA plot of the 3 different prompts with a negation prompt.
This problem has been introduced into text-to-image generation tasks, making it hard for the model to generate images without certain concepts (examples in Figure 1 of Singh et al. (2025) and Figure 5 of Park et al. (2025)). Thus, classifier-free guidance (CFG) was used to introduce a negative prompt to the image generation process. More details in the next section. Several studies have attempted to tackle this issue by employing alternative training strategies, such as incorporating harder samples in the training data designed for negation tasks (Yuksekgonul et al., 2023; Singh et al., 2024; 2025; Alhamoud et al., 2025; Park et al., 2025). Some of these methods have shown improvements in image generation tasks. For instance, Park et al. (2025) reported gains in Neg Score—measuring whether the model retains the primary subject while correctly omitting the negated object—for both SD-1.4 and SDXL-1.0, by replacing the default CLIP encoder with their NegationCLIP on their dataset, without additional T2I training.
These methods generally require re-training the text encoder (usually a CLIP-like model) with contrastive learning, which poses challenges for models that do not use contrastively pre-trained encoders, such as T5 (Raffel et al., 2023) in Stable Diffusion 3 (Stability AI, 2024; Esser et al., 2024) and Flux (Black Forest Lab, 2025). Moreover, each model using a different text encoder would require a separate, dedicated adaptation. Additionally, even if the text encoder understands the negation, the diffusion model might still fail to avoid certain items because of their strong association.
Appendix B Proof That Our Method is the Same as Token-Weighted Subtraction
In this section, we prove that our method $Z^{VSF}$ is equivalent to token-weighted subtraction, denoted $Z^{W}$.
Proof.
We define
$A^{+}=\exp(\frac{Q(K^{+})^{T}}{\sqrt{d}}),\quad A^{-}=\exp(\frac{Q(K^{-})^{T}}{\sqrt{d}}).$
Then
$W=\frac{\sum A^{+}}{\sum A^{+}+\sum A^{-}}.$
Substituting into the expression for $Z^{W}$:
$Z^{W}=W\cdot\sigma(Q(K^{+})^{T})V^{+}-(1-W)\cdot\alpha\cdot\sigma(Q(K^{-})^{T})V^{-},$
and using the softmax definitions
$\sigma(Q(K^{+})^{T})=\frac{A^{+}}{\sum A^{+}},\quad\sigma(Q(K^{-})^{T})=\frac{A^{-}}{\sum A^{-}},$
we obtain
Canceling the sums:
$Z^{W}=\frac{A^{+}}{\sum A^{+}+\sum A^{-}}V^{+}-\frac{\alpha A^{-}}{\sum A^{+}+\sum A^{-}}V^{-}.$
This matches
$Z^{VSF}=\sigma(Q(K^{+}\oplus K^{-})^{T})(V^{+}\oplus-\alpha V^{-}),$
since
$\sigma(Q(K^{+}\oplus K^{-})^{T})=\frac{A^{+}\oplus A^{-}}{\sum A^{+}+\sum A^{-}},$
and therefore
$Z^{VSF}=\frac{A^{+}}{\sum A^{+}+\sum A^{-}}V^{+}+\frac{A^{-}}{\sum A^{+}+\sum A^{-}}(-\alpha V^{-})=\frac{A^{+}}{\sum A^{+}+\sum A^{-}}V^{+}-\frac{\alpha A^{-}}{\sum A^{+}+\sum A^{-}}V^{-}.$
Hence, $Z^{W}=Z^{VSF}$. ∎
Appendix C Attention Bias and Padding Removal
In most models from Huggingface Diffusers (von Platen et al., ), padding tokens in the text input are typically not masked during attention. This is likely because the models have learned to ignore padding, and masking them would add unnecessary overhead (due to some attention implementations like FlashAttention-2 (Dao, 2023) that do not support arbitrary masking). However, when we invert the sign of the padding tokens, it degrades output quality significantly. This could be because, although these tokens carry no semantic meaning, the sign-flipping introduces unseen states into the attention mechanism. To mitigate this, we remove padding tokens from the negative prompt embeddings. For the positive prompt, we retain padding tokens, as they do not introduce novel tokens and can improve generation quality. This aligns with training conditions and may allow the model to use padding positions as registers for auxiliary information.
Appendix D Details About the Metrics
Figure 8: The left image is the original image, and the right image is generated by GPT-4o, where piano keys are removed. When scored using HPSv2, the left image got a score of 0.330 while the right image got a score of 0.319 using prompt of “A grand piano dominates an empty concert hall, a smooth ebony board stretching across the front.” and the left get a score of 0.330 and the right got a score of 0.324 if we mention “no keys” in the promps.
To evaluate the scores of the generated images, we used LLaMA 4 Maverick, which has a very high image reasoning MMMU Yue et al. (2024) score, higher than Gemma 3 and even GPT-4o. We avoided using the same model (o3) for both evaluation and generation for cost control and to avoid bias within a model. We did not evaluate the quality of the generated images using models like ImageReward (Xu et al., 2023) or HPSv2 Wu et al. (2023) as in NASA or NAG, as current quality or human preference assessment models do not account for negative prompts; traditional methods usually aim for real-world generations Ye et al. (2025). Removing a key element from the positive prompt (e.g., removing the roof from a house) is likely to reduce perceived quality, since the result deviates from what is considered “normal,” even though that is the intended outcome. An example is shown in Figure 8 where removing the keys from the piano results in a much lower quality score, even though other parts of the image look the same. Both ImageReward and HPSv2 are built on top of image-text alignment models (CLIP (Radford et al., 2021) or BLIP (Li et al., 2022)), which will likely lead to a decreased score when the main object is missing a critical part. Thus, we also let the MLLM rate the image quality from 0-1 for each image and told it to ignore the abnormality of following the negative prompt. We did observe that MLLM can make mistakes, especially when there is ambiguity or when the unwanted element is hard to see. However, we believe that in general, under 400 images, the mistakes are minor. To compensate for this, we also did a human evaluation and a more negotiation-aware MLLM evaluation to cross-validate the LLaMA evaluation. Due to model provider stability issues, we used different MLLM providers for different portions of the experiment for the same model under the same config; this could have some limited impact on the stability of metrics.
Listing 1: Pseudocode implementation of the Value Sign Flip (VSF) attention process
# prep for embeddings
pos_embeds = get_embed(prompt)
neg_embeds = get_embed(neg_prompt, padding=False)
pos_len, neg_len, img_len = pos_embeds.shape[1], neg_embeds.shape[1], IMG_LEN
# concat positive and negative prompts (N0)
prompt_embeds = torch.cat([pos_embeds, neg_embeds], dim=1)
# prep for attention mask and bias (N1 never acts as query)
total_len = img_len + prompt_embeds.shape[1]
attn_mask = torch.zeros((1, total_len, total_len + neg_len))
# block P and N0 from attending to N1
attn_mask[:,-(pos_len+neg_len):,-neg_len:] = -torch.inf
# block image and P from attending to N0
attn_mask[:,:-neg_len,-(2*neg_len):-neg_len] = -torch.inf
# block N0 and N1 from attending to P
attn_mask[:,-neg_len:,img_len:img_len+pos_len] = -torch.inf
# bias image->N1 connections
attn_mask[:,:img_len,-neg_len:] -= offset
class VSFAttnProcessor(AttnProcessor):
def __init__(self, attn_mask, neg_prompt_length):
self.attn_mask = attn_mask
self.neg_prompt_length = neg_prompt_length
def forward(self, hidden_states, encoder_hidden_states, attention_mask):
# get qkv projection for image tokens
q = self.get_q(hidden_states)
k = self.get_k(encoder_hidden_states)
v = self.get_v(encoder_hidden_states)
# get qkv projection for encoder tokens
q_enc = self.get_q_encoder(encoder_hidden_states)
k_enc = self.get_k_encoder(encoder_hidden_states)
v_enc = self.get_v_encoder(encoder_hidden_states)
query = torch.cat([q, q_enc], dim=2)
# append P, N0 (in k_enc and v_enc) and N1 (the last portion of k_enc and v_enc) at the end
k = torch.cat([k, k_enc, k_enc[:,:,-self.neg_prompt_length:]], dim=2)
v = torch.cat([v, v_enc, v_enc[:,:,-self.neg_prompt_length:]], dim=2)
# sign-flip values of N1
v[:,:,-self.neg_prompt_length:] *= -scale
hidden_states = F.scaled_dot_product_attention(
query, k, v,
dropout_p=0.0, is_causal=False,
attn_mask=self.attn_mask.to(query.dtype)
hidden_states = hidden_states.transpose(1, 2).reshape(batch_size, -1, attn.heads * head_dim)
return self.out_proj(hidden_states)
for block in model.transformer.blocks:
block.attn1.processor = VSFAttnProcessor(attn_mask, neg_len)
# diffusion process continues
Appendix E Hyper-parameter Tunning
Although NAG (Chen et al., 2025a) also targeted negative concept avoidance, its primary focus was on its effects on improving generation quality (using words like “blurry” or “low quality” as a negative prompt). We believe the hyperparameters reported in their work were tuned with an emphasis on quality rather than negation handling. Therefore, we re-tuned their hyperparameters moderately and manually on guidance scale ($\phi$), blending factor ($\alpha$), and normalization factor ($\tau$). We will report experimental results on both original NAG (noted as NAG) and the improved hyperparameter version (noted as NAG Strong). The final hyperparameters used are $\phi=11,\alpha=0.5,\tau=5$. This pushes the NAG to the edge of acceptable visual quality.
Similarly, for our VSF, we used two set of hyperparameters, VSF Quality ($\alpha=3.3,\beta=0.2$) maintained high quality and positive prompt alignment, while VSF Strong ($\alpha=3.8,\beta=0.2$) pushes it to the limit, reaching higher negative prompt alignment in trade of positive and quality score.
Appendix F Human Validation
To verify the results of the MLLM evaluation, we selected 10 prompts (with 2 seeds each) for VSF, NASA, NAG, and NAG Strong and manually labeled them with positive, negative, and quality scores. The human-labeled results are presented in Table 3. We validated MLLM performance by computing the binary F1 score and accuracy between MLLM outputs and human annotations. Cohen’s kappa was not applied due to the highly imbalanced class distribution. The reliability metrics are summarized in Table 4. We observed that quality ratings from MLLM and human labels were uncorrelated in high-quality regions. To investigate this further, we evaluated quality scores over a broader set of conditions. With a large sample size, we found that correlation emerges in a wider range: when scores are close to 1, small fluctuations carry little meaning, but substantially lower scores (e.g., <0.9<0.9) may indicate degraded quality. The correlation is shown in Figure 9. This supports the observation in Figure 11, where MLLM tends to overestimate quality. From the scatter plot and regression, we can see that the MLLM score is usually higher than the human score, and although they are not linearly correlated, they are monotonically correlated.
Figure 9: Correlation Between the Human-rated quality score and MLLM-rated quality score
Table 4: Reliability Metric For Human And MLLM Evaluation Results
Negative Score Positive Score
F1 0.667 0.974
Accuracy 0.850 0.950
Appendix G Adapting to Other DiT Models
@@
In this paper, we primarily use SD-3.5 (Stability AI, 2024) due to simplicity and elegant architecture. However, our method can theoretically be adapted to any transformer-based diffusion or flow-matching model. To demonstrate this adaptability, we implemented our method on Wan 2.1 with CausVid LoRA (Yin et al., 2025; Seppanen, ) and Flux Schnell (Black Forest Lab, 2025).
@@
For Wan 2.1, which uses cross-attention between image and text, duplication and masking are unnecessary and not used. Because our approach does not perform extrapolation and solely provides negative guidance, it cannot enhance overall quality significantly or replace CFG sampling in non-disstilled models, making it incompatible with the original Wan 2.1 model. Instead, we utilize CausVid (Yin et al., 2025), which enables Wan to function effectively without classifier-free guidance in few-step settings. Specifically, we used a LoRA distilled from the original CausVid that can be directly applied on top of Wan 2.1 (Seppanen, ). For qualitative results from Wan, please see the appendix.
We also tested our method on Flux Schnell (Black Forest Lab, 2025). However, due to the model likely being trained to associate items with their associated items that often appear together strongly, we need to make some modifications. Before the negative prompt was fed into the model, we did a CFG-like extrapolation on the negative prompt, with a mean padding embedding as a null condition. This follows the implementation of the Compel package. Noted as:
$p^{-}=p^{-}+\lambda\cdot(p^{-}-p^{\emptyset}),$ (10)
we used $\lambda=8$ in this case. Quantitative results are shown in Table 5. We can observe that even without any negative guidance, the Flux Schnell model can slightly better avoid the items solely based on the positive prompts (since the positive prompts implied the item is missing using terms like “empty”), but with the help of VSF, it further increases the negative score without compromising the positive and quality score.
Table 5: Comparsion of Flux Schnell VSF and Original Schnell (Black Forest Lab, 2025)
Method Positive Score Negative Score Quality Score
Flux Schnell 1.00 0.22 0.99
Flux Schnell VSF 0.97 0.41 0.99
Appendix H Computational Cost
@@
Since our method does not require two passes through the entire model (as in CFG) or the attention module (as in NAG or NASA), and only slightly increases the sequence length ($0.2\%$), its theoretical computational cost is significantly lower, close to that of a single pass. However, due to implementation limitations (specifically, FlashAttention-2’s lack of support for arbitrary attention masking), the actual runtime of our method is higher than the original single-pass MM-DiT models, and similar to NAG or NASA, but still lower than CFG.
@@
To accurately measure the computational cost, we evaluate the runtime of 25 identical prompts under four settings: no guidance, NAG, NASA, and our proposed guidance, VSF, and then report the average runtime and peak memory usage for each setting. We also reported the theoretical CFG time as double the one without guidance. To avoid GPU thermal throttling affecting the results, we pause for at least 5 minutes between each set of tests. The tests are done on NVIDIA A100 40GB on Google Colab, as this is the most accessible option for high-end GPUs for users. Stable-Diffusion-3.5-Large-Turbo is generated in 8 steps for 1024x1024 resolution, Wan is generated in 8 steps with 480x832 resolution, and 81 frames. The results are shown in Table 6.